given f(x)=-3x+4 and g(x)=x^2-5 find the following:
f(2) + g(2)
f(2)- g(2)
g(2)/f(2)
f(3)*g(3)

Answer by nabla(475)

Figure out f(2) and g(2), and f(3) and g(3) first.
f(2)=-3(2)+4=-2
g(2)=2^2-5=-1
f(3)=-3(3)+4=-5
g(3)=3^2-5=4
So,
f(2) + g(2)=-2-1=-3
f(2)- g(2)=-2-(-1)=-1
g(2)/f(2)=-1/(-2)=1/2
f(3)*g(3)=-5(4)=-20.